OpenStudy (anonymous):

Use implicit differentiation for the function: yz = ln(x+z) to find the partial derivative of both y and z in respect to x

OpenStudy (anonymous):

Any ideas on how to start this question?

OpenStudy (irishboy123):

for \(\dfrac{\partial z}{\partial x}\) just do normal implicit diff with \(z = z(x)\) and \(y = const\) vice versa for \(\dfrac{\partial y}{\partial x}\) that make sense?!

OpenStudy (anonymous):

yeah I got something like: \[ \frac{ ∂y}{ ∂x }z = \frac{ 1 }{ x+z }\]

OpenStudy (anonymous):

Does it seem about right? And then from here, I would make ∂y/∂x the subject

OpenStudy (irishboy123):

looks good

OpenStudy (irishboy123):

yeah. make it the subject

OpenStudy (anonymous):

Okay cool, I guess this would also be a similar outcome for ∂z/∂x

OpenStudy (irishboy123):

for z, i get \( y \dfrac{\partial z}{\partial x} = \dfrac{1}{x+z} (1 + \dfrac{\partial z}{\partial x})\) and then tidy up

OpenStudy (anonymous):

Okay cool got it!